Consider the recurrence relation an= 3 an1 + 3 an-2 when ao = 1 and a1= 2 then a3= ..: O 56


1
Expert's answer
2021-12-09T17:06:04-0500

an=3an1+3an2a_n= 3 a_{n-1} + 3 a_{n-2}


a2=3a1+3a0=6+3=9a_2=3a_1+3a_0=6+3=9


a3=3a2+3a1=39+32=33a_3=3a_2+3a_1=3\cdot9+3\cdot2=33